OpenStudy (anonymous):

Let line t be the line represented by 3x+4y = 5 and let line p be the line perpendicular to line t and containing the point (5,5). What is the x-coordinate of the point common to line t and line p?

OpenStudy (nincompoop):

First, arrange the linear equation into slope-intercept form \(\large y = mx + b \)

OpenStudy (nincompoop):

Two perpendicular lines are defined by their negative reciprocal slopes. Meaning that if the slope of the first line is, \(m = \large \frac{2}{5} \) then the slope of the other line perpendicular to it is \(m = \large - \frac{5}{2} \)

OpenStudy (anonymous):

Slope is -3/4, next

OpenStudy (nincompoop):

Upon performing the first step, that is arranging the linear equation into slope-intercept form, you will be able to determine its slope, and then determine the slope of other line with coordinates (5, 5) since it is said to be perpendicular.

OpenStudy (nincompoop):

so, if the slope is \(m \large = -\frac{3}{4} \) then what it is its negative reciprocal?

OpenStudy (anonymous):

this is the -ve reciprocal